n to the Owens Valley 1632 m below. If the 21.0 m along the road to the top of the hill. If the cyclists initial speed is 9.30 m/s, what is the final speed? Ignore friction and air
Answer #1

\\7)\\Initial\;energy\;at\;the\;top\;E_i=mgh\\\\\\ final\;energy\;at\;the\;bottom\\\\\\E_f=\frac{1}{2}mv^2\\\\\\ total\;energy\;is\;conserved\\\\\\E_f=E_i\\\\\\ v=\sqrt{2gh}=\sqrt{2\times 9.8\times 1632}=178.85\;m/s

\\8) \\Initial\;energy\;at\;the\;bottom\;E_i=\frac{1}{2}mv_i^2\\\\\\ final\;energy\;at\;the\;top\\\\\\E_f=KE_f+PE_f=\frac{1}{2}mv_f^2+mgh\\\\ h=L\;sin\theta\\\\ E_f=\frac{1}{2}mv_f^2+mgLsin\theta\\\\\\ total\;energy\;is\;conserved\\\\\\E_f=E_i\\\\\\ v_f=\sqrt{v_i^2-2gLsin\theta}=\sqrt{9.3^2-2\times 9.8\times 21\sin11.1}=2.69\;m/s\\Initial\;energy\;at\;the\;top\;E_i=mgh\\\\\\ final\;energy\;at\;the\;bottom\\\\\\E_f=\frac{1}{2}kx^2\\\\\\ total\;energy\;is\;conserved\\\\\\E_f=E_i\\\\\\ x=\sqrt{\frac{2mgh}{k}}=\sqrt{\frac{2\times 4.3\times 9.8\times 2.5}{2.38\times 10^4}}=0.094\;m=9.4\;cm

\\9)\\\\(a)\\ Potential \;energy\;stored\;in\;bow=kinetic\;energy\;of\;the\;arrow\\\\\\ PE=\frac{1}{2}mv^2=\frac{1}{2}5.25\times 10^{-2}\times 39^2=40\;J\\\\\\ (b)potential\;energy\;PE=\frac{1}{2}kx^2\\\\\\averge\;force\;F_{avg}=\frac{1}{2}kx=\frac{PE}{x}\\\\F_{avg}=\frac{40}{0.3}=133.3N